### v2

# CATSplat: Context-Aware Transformer with Spatial Guidance for Generalizable 3D Gaussian Splatting from A Single-View Image

Abstract
Recently, generalizable feed-forward methods based on 3D Gaussian Splatting have gained significant attention for their potential to reconstruct 3D scenes using finite resources. These approaches create a 3D radiance field, parameterized by per-pixel 3D Gaussian primitives, from just a few images in a single forward pass. However, unlike multi-view methods that benefit from cross-view correspondences, 3D scene reconstruction with a single-view image remains an underexplored area. In this work, we introduce CATSplat, a novel generalizable transformer-based framework designed to break through the inherent constraints in monocular settings. First, we propose leveraging textual guidance from a visual-language model to complement insufficient information from a single image. By incorporating scene-specific contextual details from text embeddings through cross-attention, we pave the way for context-aware 3D scene reconstruction beyond relying solely on visual cues. Moreover, we advocate utilizing spatial guidance from 3D point features toward comprehensive geometric understanding under single-view settings. With 3D priors, image features can capture rich structural insights for predicting 3D Gaussians without multi-view techniques. Extensive experiments on large-scale datasets demonstrate the state-of-the-art performance of CATSplat in single-view 3D scene reconstruction with high-quality novel view synthesis.
Website: https://kuai-lab.github.io/catsplat2025.
1 Introduction
3D scene reconstruction and novel view synthesis are fundamental tasks in modern computer vision and graphics, driving advancements across diverse domains [13, 29, 2, 34], such as virtual reality and autonomous navigation. Together, they create 3D scene representations using 2D source images and produce realistic images from unseen perspectives. Early approaches [35, 38, 9, 6] (e.g., NeRF) have made impressive progress through differentiable volume rendering. However, they are still far from real-time scenarios due to the heavy computational demands. Unlike previous methods, 3D Gaussian Splatting (3DGS) based approaches [22, 60, 57] have emerged as leading frontrunners, achieving high performance with real-time rendering capabilities. They employ 3D Gaussians for explicit scene representations via efficient rasterization-based rendering.
Figure 1: Overview of the generalizable 3D scene reconstruction pipeline. The feed-forward network creates a 3D radiance field using 3D Gaussians, all within an end-to-end differentiable system.
Recently, generalizable feed-forward methods [8, 10, 61, 52, 45] based on 3DGS [22] have attracted growing interest for their ability to reconstruct 3D scenes, even with constrained resources like sparse view images. They create a 3D radiance field parameterized by per-pixel Gaussian primitives from just a few input images (typically one or two) in a single forward pass without scene-specific optimization. For example, pixelSplat [8] samples Gaussian centers from a probabilistic depth distribution using a multi-view epipolar transformer, while MVSplat [10] constructs cost volumes from two source images to extract geometric cues. Both methods benefit from cross-view correspondences between a pair of images to capture useful cues for the precise prediction of Gaussian parameters. However, in contrast to the multi-view settings, which provide relatively abundant information, single-view 3D reconstruction solely depends on a single image, leading to limited cues. Although Flash3D [45] has pioneered a 3DGS-based generalizable single-view 3D scene reconstruction with a foundation monocular depth estimation model [39], this area has yet to be fully explored. Note that we outline a single-view generalizable 3D scene reconstruction pipeline in  Fig. 1.
Figure 2: We introduce CATSplat, a Context-Aware Transformer with Spatial Guidance for Generalizable 3D Gaussian Splatting from a single image. (a) Our two main priors, and (b) Examples of text descriptions (from the VLM) representing an input image.
To tackle the challenges in monocular scenarios, we introduce CATSplat, a carefully designed transformer that leverages two intelligent guidance to supplement the insufficient information from a single image. Based on the traditional paradigm of generalizable 3DGS frameworks, which predict Gaussian primitives from image features, we focus on enhancing these features with essential knowledge. First, we propose using text guidance as contextual priors. One of the most promising ways to employ text guidance is through visual-language models (VLM) [30, 27, 1, 66]. They have showcased their potential to provide visual-linguistic knowledge learned from large-scale multimodal data in various vision tasks [23, 67, 20, 24]. Motivated by the success of VLMs, we utilize text embeddings from VLM representing the input image to guide the network towards context-aware 3D scene reconstruction, as shown in Fig. 2 (a). Specifically, within cross-attention layers, we softly integrate scene-specific details of text features into image features. Here, as illustrated in Fig. 2 (b), text features encoding such descriptions can provide corresponding spatial context (e.g., kitchen) and information about objects (e.g., refrigerator and oven) usually found in these environments. These extra details can serve as valuable guidance (or bias) for effective scene reconstruction, further improving generalizability beyond relying on visual clues.
In addition to contextual guidance, we explore additional avenues to enrich the knowledge of image features. In generalizable tasks with sparse images, gaining insights into 3D geometric properties is crucial to accurately reconstruct scenes in 3D space. Typically, multi-view methods [8, 10] utilize physical techniques such as triangulation to capture comprehensive 3D cues from cross-view perspectives. However, in monocular settings, such techniques are unavailable, leading to constrained geometric details. In this context, we advocate for integrating 3D guidance into 2D features to enhance their spatial understanding. Beyond simply using a 2D depth map from an off-the-shelf depth estimation model as in previous work [45], we further leverage its 3D representation as a backprojected point cloud. As shown in Fig. 2 (a), we extract 3D features from 3D points and strengthen image features with rich structural insights of 3D features through attention mechanisms. Ultimately, our image features with two constructive priors are now highly informative for scene representation with Gaussians.
Given landmark datasets, RealEstate10K (RE10K) [65], ACID [28], KITTI [17], and NYUv2 [43], we validate the generalizability and effectiveness of our novel framework. To summarize, our main contributions are listed as follows:
  • We introduce CATSplat, a novel generalizable framework for monocular 3D scene reconstruction. We leverage the rich contextual cues of text embeddings from the VLM as insightful guidance toward context awareness, complementing limited information from a single image.
  • We propose 3D spatial guidance for a monocular image to enrich geometric details in single-view settings. With 3D priors, image features can capture valuable cues for predicting 3D Gaussians without multi-view techniques.
  • We analyze the effectiveness of our method on challenging datasets. Extensive quantitative and qualitative experiments demonstrate that ours achieves new state-of-the-art performance on single-view 3D scene reconstruction.
2 Related Work
Figure 3: Overview of CATSplat framework. CATSplat takes an image $\mathcal{I}$ and predicts 3D Gaussian primitives $\{(\bm{\mu}_{j},\bm{\alpha}_{j},\bm{\Sigma}_{j},\bm{c}_{j})\}^{J}_{j}$ to construct a scene-representative 3D radiance field in a single forward pass. In this paradigm, our primary goal is to go beyond the finite knowledge inherent in a single image with our two innovative priors. Through cross-attention layers, we enhance image features $F^{\mathcal{I}}_{i}$ to be highly informative by incorporating valuable insights: contextual cues from text features $F^{C}_{i}$, and spatial cues from 3D point features $F^{S}_{i}$.
Sparse-view 3D Reconstruction. Recent progress in neural fields [44, 55, 34] and volume rendering [31, 47] has advanced 3D reconstruction and novel view synthesis, even with sparse-view images. For example, FreeNeRF [56] regularizes frequency to address few-shot neural rendering, while pixelNeRF [58] predicts a neural radiance field in the camera coordinate using a feed-forward approach from few-view images. More recently, 3D Gaussian Splatting (3DGS) [22] has revolutionized the field of 3D reconstruction, achieving real-time rendering. Inspired by the success of 3DGS, pixelSplat [8] has pioneered the feed-forward network, which reconstructs a 3D radiance field parameterized using 3D Gaussian primitives from a pair of images. Then, diverse multi-view generalizable 3DGS approaches [10, 52, 61] have since developed with a similar structure. MVSplat [10] constructs cost volumes to capture cross-view similarities for accurate Gaussians, and latentSplat [52] introduces variational Gaussians to encode uncertainty in a latent space. While they typically benefit from cross-view properties, monocular 3D reconstruction is relatively more challenging due to limited information.
Single-view 3D Reconstruction. Early approaches [53, 49] have proposed various strategies to overcome the constraints of single-view scenarios. SynSin [53] introduces a differentiable point cloud renderer, which projects a 3D point cloud from a single image into target views. [49] predicts multiplane images (MPI) [65] directly from a single image without correlations between multiple views. In line with recent trends, single-view 3D reconstruction quality has significantly improved, thanks to innovations in NeRF [34] and 3DGS [22]. Built upon NeRF [34], MINE [25] extends MPI to a continuous 3D representation, and BTS [54] predicts less complex continuous density fields from an image. Recently, Splatter Image [46] involves 3D Gaussians for monocular object reconstruction through an image-to-image neural network. Also, Flash3D [45] predicts pixel-wise Gaussian parameters in a single forward pass without expensive per-scene optimization, relying on a foundation monocular depth estimation model [39]. Based on the core idea of the generalizable 3DGS framework, our novel approach, CATSplat, leverages two beneficial guidance to complement insufficient details from a single image.
Vision-Language Models for Vision Tasks. Visual Language Models (VLMs) have emerged as powerful tools for bridging the gap between visual and textual modalities [16, 32], achieving outstanding performance in diverse vision tasks, such as image captioning [26, 27, 3, 59, 37], image-text retrieval [42, 21, 33, 64], and visual question answering (VQA) [36, 41, 19]. These models use large-scale image-text pair datasets to learn joint representations, encouraging seamless understanding and integration across both modalities. Early approaches like CLIP [42] and ALIGN [21] leverage contrastive learning to relate image and text data within a shared embedding space, enabling effective zero-shot generalization across modalities. Recently, the success of Large Language Models (LLMs) [4, 48, 11, 7] has driven significant advancements in visual-language processing. For example, BLIP-2 [27] and LLaVA [30] demonstrate strong performance in image captioning with context-rich visual descriptions based on LLMs [11, 63, 12]. Specifically, they aim to connect image features from a visual encoder into the language space of pre-trained LLMs. In this work, motivated by the effectiveness of VLMs, we employ contextual clues of text embeddings from VLM to complement the limited information from a monocular image.
3 Method
In this section, we introduce CATSplat, a novel generalizable framework for monocular 3D scene reconstruction with 3D Gaussian Splatting. We first provide an overview of the whole pipeline (Sec. 3.1 and Fig. 3) and then elaborate on technical details: Context-Aware 3D Reconstruction (Sec. 3.2) and Spatial Guidance for 3D Insights (Sec. 3.3).
3.1 Overview
Recent generalizable feed-forward frameworks [8, 10, 61, 52, 45] commonly follow a similar paradigm; they construct a 3D radiance field from $N$ sparse-view images $\mathcal{I}^{N}\in\mathbb{R}^{N\times H\times W\times 3}$ in a single forward pass with pixel-aligned $J$ Gaussian primitives $\{(\bm{\mu}_{j},\bm{\alpha}_{j},\bm{\Sigma}_{j},\bm{c}_{j})\}^{J}_{j}$, including position $\bm{\mu}_{j}$, opacity $\bm{\alpha}_{j}$, covariance $\bm{\Sigma}_{j}$, and spherical harmonics coefficients $\bm{c}_{j}$. In this paradigm, it is challenging to reconstruct the vivid scene from a single image due to limited resources, comparing with multi-view configurations. To overcome this constraint, we propose a carefully designed transformer that leverages two extra guidance for enhancing knowledge of single-view image features: (1) Text Guidance, which provides deep contextual clues for the scene, and (2) Spatial Guidance, which enriches three-dimensional structural information of 2D features, as illustrated in Fig. 3.
Feed-Forward Network with Transformer. From a single input image $\mathcal{I}\in\mathbb{R}^{H\times W\times 3}$, we first predict a depth map $D\in\mathbb{R}_{+}^{H\times W\times 1}$ as potential centers for Gaussians, employing a pre-trained monocular depth estimation model [39]. Given $\mathcal{I}$ and its estimated depth map $D$, we channel-wise concatenate them as ${\mathcal{I}}^{\prime}\in\mathbb{R}^{H\times W\times 4}$, then feed $\mathcal{I}^{\prime}$ into a ResNet-based image encoder [18] to produce hierarchical depth-conditioned image features $F_{i}^{\mathcal{I}}\in\mathbb{R}^{H_{i}\times W_{i}\times D_{i}^{\mathcal{I}}}$. Then, we utilize a multi-resolution transformer that encourages image features $F_{i}^{\mathcal{I}}$ to effectively represent both global structures and fine details across various resolutions, improving the overall understanding of the scene. We specifically use three layers with three resolution features. Based on transformer architecture, we extend the cross-attention mechanism to interact with our two novel priors, as described in Sec. 3.2 and Sec. 3.3, further enriching the feature representation. Through iterative layers, our transformer yields highly informative image features $\tilde{F}_{i}^{\mathcal{I}}\in\mathbb{R}^{H_{i}\times W_{i}\times D_{i}^{\mathcal{I}}}$ well-suited for effective scene reconstruction in 3D space. We ultimately estimate the parameters of Gaussians from $\tilde{F}_{i}^{\mathcal{I}}$ using ResNet-based decoders, as detailed in Sec 3.4.
3.2 Context-Aware 3D Reconstruction
In real-world scenarios, diverse objects are usually placed in inconsistent patterns without following conventional rules. These complexities make monocular 3D scene reconstruction more challenging, as it depends on insufficient details available from an image. To transcend the limits of finite knowledge, we advocate leveraging textual information as a rich source of hidden context, enhancing generalizability.
Incorporation of Textual Cues. Recent advancements in large-scale visual language models [30, 27, 1, 66] (VLM) have highlighted the benefits of their general embedded knowledge, which mirrors the diversity of real-world contexts. In this work, we take advantage of generous contextual cues inherent in the text representations produced by these models. With a single-view source image $\mathcal{I}$, we prompt the pre-trained VLM [30] to generate a detailed, one-sentence description of the scene. During this procedure, we utilize text embeddings $F^{C}\in\mathbb{R}^{N_{c}\times D^{C}}$ from a well-aligned multimodal space before they are processed into linguistic descriptions. Our main focus is on the contextual details from $F^{C}$, such as object identities, spatial relationships, and scene semantics, which can potentially serve as influential biases for enhancing generalizability. To softly incorporate supplemental cues from $F^{C}$ into image features $F^{\mathcal{I}}$, we employ iterative cross-attention layers. For each transformer layer designed to use multi-scale features, we convert $F^{C}$ into $F_{i}^{C}\in\mathbb{R}^{N_{c}\times D^{C}_{i}}$ to align the dimension with its corresponding $F_{i}^{\mathcal{I}}$ using a linear layer, as illustrated in Fig. 4. Given $F_{i}^{\mathcal{I}}$ and $F_{i}^{C}$, queries $\mathbf{Q}_{i}$ are projected from $F_{i}^{\mathcal{I}}$, and keys $\mathbf{K}_{i}$ and values $\mathbf{V}_{i}$ are from $F_{i}^{C}$, as follows:
$\mathbf{Q}_{i}=W_{q}\cdot F_{i}^{\mathcal{I}},\;\>\mathbf{K}_{i}=W_{k}\cdot F_{i}^{C},\;\>\mathbf{V}_{i}=W_{v}\cdot F_{i}^{C}$ (1)
where $W$ denotes the learnable parameters of each projection layer. Then, we associate them through cross-attention:
$F_{i}^{\mathcal{I}C}=Attn(\mathbf{Q}_{i},\mathbf{K}_{i},\mathbf{V}_{i})=\text{Softmax}(\frac{\mathbf{Q}_{i}\cdot\mathbf{K}_{i}^{T}}{\sqrt{D_{i}}})\mathbf{V}_{i}\vspace{-.5em}$ (2)
where $F_{i}^{\mathcal{I}C}$ represents output features containing not only visual clues from $F_{i}^{\mathcal{I}}$ but also textual clues from $F_{i}^{C}$. Finally, our iterative layers continuously establish valuable connections between an input monocular image and additional contextual priors, facilitating more generalizable 3D reconstruction of complex scenes under limited resources.
Figure 4: Detailed transformer pipeline. In the $i$-th layer, we first operate cross-attention between $F_{i}^{\mathcal{I}}$ and $F_{i}^{C}$, then proceed cross-attention with $F_{i}^{S}$. We also use a ratio $\gamma$ to preserve visual information from $F_{i}^{\mathcal{I}}$ while incorporating extra cues from $F_{i}^{C}$ and $F_{i}^{S}$.
3.3 Spatial Guidance for 3D Insights
In multi-view configurations, each perspective contributes unique spatial information, boosting the reconstruction of complex three-dimensional structures. Yet, single-view often falls short of 3D cues for comprehensive geometric understanding. To bridge this gap, we introduce efficient spatial guidance based on the 3D representation of a 2D depth map, which provides a broader geometric context for reliable 3D perception independent of stereo vision expertise.
Incorporation of Spatial Cues. Solid geometric awareness is essential for accurately depicting a scene within 3D space. To capture 3D cues from a single image, traditional approaches [45, 46, 25] often rely on depth information in a two-dimensional format. Beyond its conventional use, we extend the estimated per-pixel 2D depth $d\in D$ into a full 3D representation for more direct spatial knowledge. Given camera parameters $K$ = $diag(f_{x},f_{y},1)\in\mathbb{R}^{3\times 3}$, where $f$ denotes the focal length, we unproject $D$ into 3D space as point cloud $P\in\mathbb{R}^{H\times W\times 3}$, with each point $\bm{p}\in P$:
$\bm{p}=K^{-1}\cdot\bm{u}\cdot d=(u_{x}d/f_{x},u_{y}d/f_{y},d)\vspace{-.2em}$ (3)
where $\bm{u}=(u_{x},u_{y},1)\in\mathcal{I}$ is one of the image pixels. From this set of points $P$, we extract 3D features $F^{S}\in\mathbb{R}^{N_{s}\times D^{S}}$ using a PointNet-based encoder [40] for better spatial reasoning. These 3D embeddings usually encode important geometric details, from depth relationships to surface orientations, going beyond static depth information. In order to integrate such valuable clues into image features while overcoming the domain gap between 2D and 3D representations, we leverage cross-attention layers. Similar to the approach for textual cues, we project $F^{S}$ into $F_{i}^{S}\in\mathbb{R}^{N_{s}\times D^{S}_{i}}$ and further enrich context-guided image features $F^{\mathcal{I}C}_{i}$ (Eq. 2) from the previous cross-attention layer with $F_{i}^{S}$ as follows:
$F_{i}^{\mathcal{I}CS}=Attn(\mathbf{Q}^{\prime}_{i},\mathbf{K}^{\prime}_{i},\mathbf{V}^{\prime}_{i})=\text{Softmax}(\frac{\mathbf{Q}^{\prime}_{i}\cdot{\mathbf{K}^{\prime}_{i}}^{T}}{\sqrt{D_{i}}})\mathbf{V}^{\prime}_{i}\vspace{-.3em}$ (4)
where $\mathbf{Q}^{\prime}_{i}$ are projected from $F_{i}^{\mathcal{I}C}$, and $\mathbf{K}^{\prime}_{i}$ and $\mathbf{V}^{\prime}_{i}$ are from $F_{i}^{S}$. During the add and normalization process after cross-attention, as shown in Fig. 4 below, we use the ratio $\gamma$ to preserve core visual information from the source image while incorporating practical cues from our two novel priors as:
$\tilde{F}_{i}^{\mathcal{I}CS}=\text{Norm}(F_{i}^{\mathcal{I}}+\bm{\gamma}\>\text{Dropout}(F_{i}^{\mathcal{I}CS}))\vspace{-.2em}$ (5)
Then, we refine $\tilde{F}_{i}^{\mathcal{I}CS}$ to $\tilde{F}_{i}^{\mathcal{I}}$ with the self-attention layer, ensuring seamless knowledge enhancement across the feature space. Ultimately, the final output features $\tilde{F}_{i}^{\mathcal{I}}$ from the transformer are now highly informative for robust scene reconstruction in tough 3D space, even with a single image.
3.4 Gaussian Parameters Prediction
With insightful features $\tilde{F}_{i}^{\mathcal{I}}$, we predict parameters for $J$ pixel-aligned 3D Gaussians $\{(\bm{\mu}_{j},\bm{\alpha}_{j},\bm{\Sigma}_{j},\bm{c}_{j})\}^{J}_{j}$ through ResNet-based decoders [18] to represent the 3D scene.
Gaussian center $\bm{\mu}$. For precise scene reconstruction, we predict depth offsets $\delta\in\mathbb{R}_{+}^{H\times W\times 1}$ to refine per-pixel depth $d\in D$ and 3D offsets $\Delta_{j}\in\mathbb{R}^{3}$ for center-wise alignment following [46, 45]. Then, we unproject the 2D refined depth $\tilde{d}=d+\delta$ into 3D points using the provided camera parameters K to produce potential centers. Given $\Delta_{j}$ and projected points, the $j^{th}$ Gaussian center $\bm{\mu}_{j}$ is set as follows:
$\bm{\mu}_{j}$ $=K^{-1}\cdot\bm{u}\cdot\tilde{d}+\Delta_{j}$ (6)
$=(u_{x}\tilde{d}/f_{x}+\Delta_{x},\;\>u_{y}\tilde{d}/f_{y}+\Delta_{y},\;\>\tilde{d}+\Delta_{z})\vspace{-.4em}$ (7)
where $\bm{u}=(u_{x},u_{y},1)\in\mathcal{I}$ is one of the image pixels.
Opacity $\bm{\alpha}$, Covariance $\bm{\Sigma}$, and Color $\bm{c}$. In line with previous generalizable feed-forward methods [8, 10] using 3DGS, we operate convolutional layers to predict each parameter. We use the sigmoid activation function for the opacity $\bm{\alpha}$ to ensure that values are bounded between 0 and 1. Additionally, we estimate a rotation matrix $R$ and a scaling matrix $S$ to construct the covariance matrix $\bm{\Sigma}=RSS^{T}R^{T}$. Also, for the color, we decode spherical harmonics coefficients $\bm{c}$.
Loss Function. Finally, we render images ^$\hat{\mathcal{I}}_{t}$ from novel viewpoints based on the reconstructed 3D scene using rasterization operation. For training, we calculate the following loss $\mathcal{L}_{total}$ as the sum of the three losses to optimize the quality of the rendered images ^$\hat{\mathcal{I}}_{t}$ with GT target images $\mathcal{I}_{t}$:
$\mathcal{L}_{total}=\lambda_{\ell 1}\mathcal{L}_{\ell 1}+\lambda_{\text{ssim}}\mathcal{L}_{\textrm{ssim}}+\lambda_{\text{lpips}}\mathcal{L}_{\textrm{lpips}}$ (8)
where $\mathcal{L}_{\textrm{ssim}}$ and $\mathcal{L}_{\textrm{lpips}}$ represent Structural Similarity Index (SSIM) and Learned Perceptual Image Patch Similarity (LPIPS) [62] losses, respectively, and each $\lambda$ is a hyper-parameter to handle the strength of the respective loss term.
4 Experiments
4.1 Experimental Setup
Datasets. In this study, we train and evaluate the overall performance using a large-scale dataset, RealEstate10K (RE10K) [65], containing home walkthrough videos. We also use three additional datasets, NYUv2 (indoor) [43], ACID (nature) [28], and KITTI (driving) [17], for cross-dataset experiments. Detailed descriptions of datasets and implementation details are provided in the supplementary.
Evaluation Metrics. We quantitatively evaluate the 3D reconstruction performance using three traditional metrics for novel view synthesis: PSNR, SSIM [51], and LPIPS [62]. For comparison with single-view 3D reconstruction methods, we evaluate three metrics on unseen target frames located 5 and 10 frames away from the input source image as well as a randomly sampled frame within a ±30 frame range, following the standard evaluation protocol of previous methods [25, 45]. Also, to further evaluate our method, we adopt conventional interpolation and extrapolation protocols from pixelSplat [8] and latentSplat [52], respectively, following Flash3D [45]. For extrapolation, we sample target views up to 45 frames before or after the source frame.
{tabu}
lccccccccccc $n=5$ (frames) $n=10$ (frames) $n=\textit{Random}$ (frames)
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
MPI [49] 27.10 0.870 – 24.40 0.812 – 23.52 0.785 –
BTS [54] – – – – – – 24.00 0.755 0.194
Splatter Image [46] 28.15 0.894 0.110 25.34 0.842 0.144 24.15 0.810 0.177
MINE [25] 28.45 0.897 0.111 25.89 0.850 0.150 24.75 0.820 0.179
Flash3D [45] 28.46 0.899 0.100 25.94 0.857 0.133 24.93 0.833 0.160
CATSplat (Ours) 29.09 0.907 0.094 26.44 0.866 0.125 25.45 0.841 0.151
Table 1: Comparisons of Novel View Synthesis (NVS) performance with state-of-the-art single-view 3D reconstruction approaches on the RealEstate10K [65] dataset. Following the standard protocol from [25, 45], we evaluate NVS metrics on unseen target frames located $n$ frames away from the input source frame. Also, we randomly sample an extra target frame within 30 frames apart from the source frame.
{tabu}
lccccccccc RE10K Interpolation RE10K Extrapolation
Input Method Framework PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Two-View
pixelNeRF [58] NeRF 20.51 0.592 0.550 20.05 0.575 0.567
Du et al. [14] NeRF 24.78 0.820 0.213 21.83 0.790 0.242
pixelSplat [8] 3DGS 26.09 0.864 0.136 21.84 0.777 0.216
latentSplat [52] 3DGS 23.93 0.812 0.164 22.62 0.777 0.196
MVSplat [10] 3DGS 26.39 0.869 0.128 23.04 0.813 0.185
Single-View
Flash3D [45] 3DGS 23.87 0.811 0.185 24.10 0.815 0.185
CATSplat (Ours) 3DGS 25.23 0.835 0.159 25.35 0.837 0.159
Table 2: Comparisons of NVS performance with state-of-the-art few-view 3D reconstruction approaches on the RealEstate10K [65]. Although we mainly focus on comparing with the leading single-view method, Flash3D [45], we also provide scores of two-view methods for additional references. Following Flash3D, we use interpolation and extrapolation protocols from previous works, [8] and [52], respectively.
4.2 Performance Comparison with SOTA Methods
Comparison with Single-view Methods. In this section, we quantitatively compare our proposed framework CATSplat with existing state-of-the-art single-view 3D reconstruction methods [53, 49, 54, 46, 25, 45]. Despite significant advancements through robust radiance field rendering techniques [34, 22], monocular 3D scene reconstruction has yet to be fully explored and still faces challenges under resource constraints. To address this challenging task, we introduce a carefully designed transformer-based architecture with two novel priors, enriching image features to predict precise 3D Gaussians for scene representation. As reported in Tab. 4.1, we evaluate novel view synthesis performance on the RealEstate10K [65] dataset. CATSplat consistently outperforms previous methods with new state-of-the-art scores in terms of PSNR, SSIM, and LPIPS across three target frame at distinct locations. Specifically, CATSplat achieves high-quality rendering not only for nearby frames, such as those 5 or 10 frames apart, but also for frames randomly located at far distances (within a ±30 frame range). These results demonstrate that our proposed priors effectively complement limited information available from a single image.
Interpolation and Extrapolation. In multi-view setups, novel view synthesis is typically evaluated on target frames within the range of multiple input images (interpolation) and outside their range (extrapolation). In Tab. 4.1, to further validate our method, we evaluate CATSplat across both conventional settings, as established in Flash3D [45], a prominent single-view 3D reconstruction method. While our primary focus is on comparing with Flash3D, we also provide scores of multi-view methods [58, 14, 8, 52, 10] for additional references. First, CATSplat significantly surpasses Flash3D in the interpolation setup. Although our results are somewhat lower than recent two-view methods, which are robust for intermediate views via cross-view correspondence, ours achieves competitive scores. Moreover, for extrapolation, CATSplat outperforms Flash3D by large margins. Notably, these impressive scores even exceed previous two-view methods despite using only a single image. In such extrapolation, target frames are usually over 45 frames away from the source image, representing nearly unseen views. These findings confirm the efficacy of our novel priors, providing helpful insights for handling distant target views. Specifically, contextual cues from text features, such as object identities (e.g., sofa, table) and scene semantics (e.g., living room), alongside spatial cues from 3D features, such as depth relationships, effectively enhance generalizability, even in challenging settings with sparse information.
Cross Dataset Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
RE10K $\rightarrow$ NYU Flash3D [45] 25.09 0.775 0.182
CATSplat (Ours) 25.57 0.781 0.157
RE10K $\rightarrow$ ACID Flash3D [45] 24.28 0.730 0.263
CATSplat (Ours) 24.73 0.739 0.250
RE10K $\rightarrow$ KITTI Flash3D [45] 21.96 0.826 0.132
CATSplat (Ours) 22.43 0.833 0.122
Table 3: Comparisons of cross-dataset generalization with the state-of-the-art single-view 3DGS method, Flash3D [45], on various real-world datasets: NYU [43], ACID [28], and KITTI [17].
{tabu}
ccccccccccccccc Method $n=\text{5}$ (frames) $n=\text{10}$ (frames) $n=\textit{Random}$ (frames)
Baseline Contextual Spatial   PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
$\checkmark$ - - 28.61 0.900 0.099 26.04 0.857 0.132 25.02 0.834 0.159
$\checkmark$ $\checkmark$ - 29.04 0.904 0.097 26.40 0.864 0.127 25.40 0.838 0.153
$\checkmark$ - $\checkmark$ 29.03 0.905 0.095 26.38 0.864 0.127 25.42 0.837 0.153
$\checkmark$ $\checkmark$ $\checkmark$ 29.09 0.907 0.094 26.44 0.866 0.125 25.45 0.841 0.151
Table 4: Ablation study to investigate the effect of our two intelligent priors (Contextual and Spatial) across three different settings, as in Tab. 4.1, on the RealEstate10K [65] dataset. Here, the “Baseline” indicates our basic transformer architecture without any proposed priors.
Figure 5: Ablation study to see the effect of iteratively incorporating our novel priors on the RE10K [65] ($n$=Random). For clear ablations, we keep the number of entire transformer layers consistent across the experiments and adjust only the number of cross-attentions (CA).
$n=\text{10}$ (frames) $n=\textit{Random}$ (frames)
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Baseline 26.04 0.857 0.132 25.02 0.834 0.159
w/ Scene Type 26.14 0.859 0.130 25.13 0.835 0.158
w/ Object List 26.23 0.862 0.128 25.25 0.836 0.155
w/ Extended 26.31 0.862 0.128 25.29 0.837 0.154
w/ Single Sent. 26.40 0.864 0.127 25.40 0.838 0.153
Table 5: Ablation study to see the impact of various text description formats for contextual guidance, including Scene Type (e.g., kitchen), Object List (e.g., oven, stove), Single Sentence, and Extended Sentences (more than two). The “Baseline” is as in Tab. 4.2.
Cross-dataset Generalization. In Tab. 3, we demonstrate the strong generalizability of CATSplat across three different cross-dataset settings. In each case, we train our model on RE10K [65] and directly test it on the target datasets in a zero-shot manner. We first evaluate the generalization on the NYU [43], which contains indoor scenes similar to the RE10K. CATSplat adeptly synthesizes images for previously unseen indoor environments. Then, we focus on outdoor scenarios with more significant domain gaps; specifically, the ACID [28] includes nature landscapes captured by aerial drones, and KITTI [17] comprises driving scenes tailored for autonomous driving. Within these challenging conditions, where filming techniques (e.g., drone) or object types (e.g., cars, buildings) are dissimilar, CATSplat showcases superior generalizability than the latest method, Flash3D [45]. Through a series of rigorous experiments, we prove the power of our intelligent priors, which empower informativeness for generalizable 3D reconstruction across real-world scenes beyond the finite scope of a single image.
Figure 6: Qualitative comparisons of NVS performance between Flash3D [45] and ours with Ground Truth on the novel view frames from RealEstate10K [65] and ACID [28] (cross-dataset). We provide more visual results and details of user study in the supplementary material.
4.3 Ablation Studies
Effect of Contextual and Spatial Priors. In Tab. 4.2, we evaluate variants of our method with/ and w/o Contextual and Spatial priors. Here, the Baseline refers to our basic multi-resolution transformer architecture, excluding cross-attention with any of our proposed priors. The addition of each prior consistently enhances the visual quality of the rendered images from target novel perspectives. With contextual priors, the improvements across all metrics underscore the significance of incorporating extra context details for effective scene reconstruction. Also, spatial priors contribute impressive gains within all target settings, providing a more extensive geometric context for rich 3D understanding. Ultimately, combining both valuable priors together leads to further advancements, achieving the best scores. These results highlight that each prior plays a meaningful role in complementing limited cues from a single image.
$n=\text{10}$ (frames) $n=\textit{Random}$ (frames)
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Baseline 26.04 0.857 0.132 25.02 0.834 0.159
w/o Depth Conc. 25.91 0.855 0.134 24.82 0.827 0.165
w/ Point Conc. 26.06 0.857 0.132 25.04 0.834 0.158
w/ Depth Feat. 26.18 0.859 0.130 25.16 0.835 0.157
w/ Point Feat. 26.38 0.864 0.127 25.42 0.837 0.153
Table 6: Ablation study to explore strategies for enhancing geometric knowledge from a single image. Here, Conc. denotes concatenation, and Feat. is features. The “Baseline” is as in Sec. 4.2.
Iteratively Incorporating Priors. Based on transformer, our feed-forward network seamlessly integrates insights from two additional priors via iterative cross-attention layers. In Fig. 5, we explore the effect of varying the cross-attention iterations using rendered images with corresponding error maps. Specifically, we keep the total transformer layers consistent at three and apply cross-attention either in the first layer only, across two layers, or throughout all three layers. Across experiments, increasing iteration of cross-attention leads to more precise, less blurry image synthesis with fewer errors. These improvements in visual quality through iterative incorporation underline the potential of our priors, providing valuable cues for 3D reconstruction.
Analysis of Context Details. We prompt a well-trained VLM [30] to generate a text description representing an input image; then, we utilize intermediate text embeddings. Here, we investigate how various context details embedded in these text features influence generalizability. In Tab. 5, we conduct experiments with four different prompt styles: identifying the scene type (e.g., bedroom), listing objects (e.g., lamp, bed), describing the scene with a detailed single sentence, and two or more sentences. While scene type or object list offers certain clues, their impact on performance is relatively modest. In contrast, sentence-level text embeddings contain more practical context details, such as texture, object relationships, and overall composition, for enhancing generalizability. However, overly extended versions may include overstatements. We ultimately employ single-sentence embeddings that provide proper details yet unexaggerated context knowledge, performing optimal scene reconstruction. We further discuss text descriptions in Supp.
Analysis of Geometric Cues. To capture geometric cues under limited resources, it is crucial to guide the network with practical spatial information. In Tab. 6, we examine strategies to enrich geometrical knowledge from a single image. Our base transformer network, called Baseline, concatenates depths with an image to extract depth-conditioned features. We first evaluate using only the image, excluding depth concatenation, and observe drops in overall scores. This highlights the meaningful role of the geometric condition. Then, we replace the depth concatenation in the Baseline with unprojected 3D point concatenation. While using 3D points yields slight gains, there is no significant benefit over depth. Beyond simple concatenation, we employ attention strategies to integrate geometric cues seamlessly. We finally observe that cross-attention with 3D point features greatly contributes to comprehensive 3D understanding, achieving potent scores than 2D depth features. These validate the efficacy of our spatial guidance incorporation.
4.4 Visual Comparison
Qualitative Analysis. In Fig. 6, we qualitatively compare rendered images from ours and Flash3D [45], along with ground truth for solid comparisons. In Scene 1 (chair) and 2 (sink), ours achieves more precise object placement with less blurriness compared to previous work. Also, in Scene 3 (stair), CATSplat clearly represents a low-texture area, whereas Flash3D struggles with blotchy artifacts. Moreover, ours outperforms Flash3D in cross-dataset scenarios. In Scene 4 and 5, our method captures well-defined edges; also, in Scene 6, ours renders a more detailed image from an aerial view of the complex cityscape. In addition to comparing rendered RGBs, we qualitatively assess the quality of 3D Gaussians for scene representation. In Fig. 7, ours predicts clearer Gaussians than Flash3D, which exhibits messy artifacts. Our excellence is also evident in the depth maps produced by these Gaussians. These findings confirm our two priors boost monocular 3D reconstruction performance.
User Study. In Tab. 7, we validate our method through human evaluation. We randomly selected 60 and 20 scenes from the RE10K [65] and ACID [28] datasets, and recruited 100 participants via Amazon Mechanical Turk. We present two types of questions with rendered images: (i) preferring between ours and Flash3D [45] based on performance, and (ii) rating the visual quality on a 7-point Likert scale. For all evaluations, ours strongly outperforms Flash3D by a significant margin across both datasets. Also, the narrow confidence interval highlights the consistency of these results.
Figure 7: Qualitative comparisons of 3D reconstruction between Flash3D [45] and ours with Ground Truth. We visualize zoom-in views of 3D Gaussians and depth maps from these Gaussians.
RE10K [65] ACID [28]
[…]
Flash3D [45] 11.58$\pm$1.09 4.56$\pm$0.30 8.59$\pm$0.63 4.14$\pm$0.21
CATSplat (Ours) 88.42$\pm$1.09 6.04$\pm$0.22 91.41$\pm$0.63 5.27$\pm$0.18
Table 7: User study comparisons. We report mean preference percentage and a 7-point Likert scale with a 95% confidence interval.
5 Conclusion
We introduce CATSplat, a novel generalizable 3DGS frame work using a single-view image. Our core objective is to transcend the constraints of relying on a single image. To this end, we propose two priors: (i) contextual priors from VLM text embeddings towards context-aware 3D scene reconstruction, and (ii) spatial priors from 3D point features for comprehensive geometric understanding. Extensive experiments demonstrate the superiority of CATSplat. While our method excels in monocular 3D scene reconstruction, ours might be less effective in occluded or truncated areas. Besides, our current training relies on the RealEstate10K dataset; however, with diverse large-scale datasets, CATSplat would be more suitable for real-world applications.
Supplementary Material
Overview
In this supplementary material, we provide further explanations and visualizations of our main paper, “CATSplat: Context-Aware Transformer with Spatial Guidance for Generalizable 3D Gaussian Splatting from A Single-View Image”. First, we elaborate on the specifics of our user study (Sec. 6). Then, we present additional technical details on the CATSplat architecture (Sec. 7). Also, we describe the implementation and datasets in more detail (Sec. 8). Moreover, we provide more quantitative and qualitative experimental results to further validate the robustness of CATSplat for 3D reconstruction and novel view synthesis (Sec. 9). Finally, we discuss the limitations of our approach (Sec. 10).
6 User Study Details
We conduct a user study to validate our method from the perspective of human perception, as described in Sec.4.4 in the main paper. Through Amazon Mechanical Turk (AMT), a widely used platform for user studies, we recruited 100 participants. We randomly sample 60 scenes from the RE10K [65] evaluation set and 20 from the ACID [28] evaluation set. Then, we use rendered images from sampled scenes for the survey questions. With rendered images and corresponding ground truth target images, we ask two types of questions, as shown in Fig. 8. For the first type of question, we show two rendered images, one from CATSplat and the other from Flash3D [45], along with a target image, and ask, “Which of the two images predicts the target image better in terms of visual quality, such as object appearance, shapes, colors, and textures?”. For the second type of question, we request participants to rate the visual quality of the rendered image from either method (CATSplat or Flash3D) on a 7-point Likert scale, with the question, “How good is the quality of the rendered image compared to the target image?”. We also include control questions to verify the reliability of responses from each participant by displaying the ground truth image as the rendered image and asking participants to rate it based on the same ground truth image, where the results are expected to be obviously high. Moreover, the method names are anonymized and presented in random order to minimize bias. We finally gathered 9,000 responses on RE10K and 6,000 responses on ACID (i.e., 30 questions for type one and 30 rating questions for each CATSplat and Flash3D on RE10K, as well as 20 questions for type one and 20 rating questions for each on ACID). Given responses from all participants, we report scores with 95% confidence intervals, as shown in Tab.7 of the main paper. Specifically, for the first type of question, which requires participants to choose between two rendered images, we utilize a binomial proportion confidence interval to analyze preferences. In the case of the second type, which queries to rate the visual quality of a single rendered image, we use a normal distribution confidence interval to analyze the average rating score. Ultimately, the results underscore the superiority of our method, as CATSplat is notably preferred and receives higher ratings compared to the latest method.
Figure 8: Examples of two types of user study questions. The first type of question (above) asks about preference between ours and Flash3D [45], and the second (below) requires participants to rate the visual quality of the rendered image compared to the target.
7 Architecture Details
Figure 9: Detailed architecture of 3D point feature extraction from a monocular input image $\mathcal{I}$. Our point cloud encoder takes back-projected points $P$ and produces point features $F^{S}$ based on the PointNet [40] structure. Here, T-Net denotes an affine transform network.
Figure 10: Examples of input images with their corresponding estimated depth maps and back-projected 3D point clouds. For better visualization, we also show 3D point clouds with RGB colors.
7.1 Details on 3D Point Feature Extraction
As described in Sec 3.3 in the main paper, we advocate incorporating 3D priors from 3D point features, which contain more comprehensive 3D domain knowledge than 2D depth maps, to address limited geometric information inherent in single-view settings. In this section, we provide additional explanations on the procedure of producing 3D point features from a single source image. As illustrated in Fig. 9, our approach first extracts a pixel-wise depth map $D\in\mathbb{R}_{+}^{H\times W\times 1}$ from an input image $\mathcal{I}\in\mathbb{R}^{H\times W\times 3}$ using a pre-trained monocular depth estimation model [39]. Next, we back-project $D$ into a 3D point cloud $P\in\mathbb{R}^{H\times W\times 3}$ with the corresponding camera parameters $K\in\mathbb{R}^{3\times 3}$. Then, a point cloud encoder takes $P$ to yield point features $F^{S}\in\mathbb{R}^{N_{s}\times D^{S}}$. Here, we organize our point cloud encoder based on the prevalent PointNet [40] architecture. Given the points $P$, we sample $N_{s}$ points using the Farthest Point Sampling (FPS) [15] algorithm; then, these sampled points $P^{\prime}\in\mathbb{R}^{N_{s}\times 3}$ are processed through a series of joint alignment networks and MLP layers. The first alignment network maps the sampled points $P^{\prime}$ to a canonical space, and the second aligns intermediate features ${F^{\prime S}}\in\mathbb{R}^{N_{s}\times 64}$ to a joint feature space. Both networks employ an affine transform matrix predicted by the T-Net. Finally, we produce 3D point features $F^{S}\in\mathbb{R}^{N_{s}\times D^{S}}$, where $D^{S}$ denotes 1,024. In Fig. 10, we present examples of input images $\mathcal{I}$, along with their corresponding depth maps $D$ and back-projected 3D point clouds $P$ (+ w/ RGB), to help understand our process.
7.2 CATSplat Procedure
In Algorithm. 1, we present the overall workflow of our generalizable feed-forward network, incorporating two novel priors, for 3D scene reconstruction from a single image.
Input: A monocular image $\mathcal{I}\in\mathbb{R}^{H\times W\times 3}$
Result: Novel view images ^$\hat{\mathcal{I}}_{t}\in\mathbb{R}^{H\times W\times 3}$
Procedure:
1
2Estimate Depth Map $D$ from $\mathcal{I}$.
3Concatenate $\mathcal{I}$ and $D$ as $\mathcal{I}^{\prime}$.
4Extract multi-resolution image features $F_{i}^{I}$ from $\mathcal{I}^{\prime}$.
5Produce text features $F_{i}^{C}$ based on the VLM.
6Back project $D$ into 3D points $P$.
7Produce 3D point features $F_{i}^{S}$ from $P$.
# Multi-resolution Transformer with $N_{l}$ layers.
8
9for $i=1$ to $N_{l}$ do
10       # Incorporation of Contextual Cues.
11      $\mathbf{Q}_{i},\mathbf{K}_{i},\mathbf{V}_{i}=W_{q}\cdot F_{i}^{\mathcal{I}},\;\>W_{k}\cdot F_{i}^{C},\;\>W_{v}\cdot F_{i}^{C}$
12      $F_{i}^{\mathcal{I}C}=Attn(\mathbf{Q}_{i},\mathbf{K}_{i},\mathbf{V}_{i})$
13      # Incorporation of Spatial Cues.
14      $\mathbf{Q}^{\prime}_{i},\mathbf{K}^{\prime}_{i},\mathbf{V}^{\prime}_{i}=W_{q}^{\prime}\cdot F_{i}^{\mathcal{I}C},\;\>W_{k}^{\prime}\cdot F_{i}^{S},\;\>W_{v}^{\prime}\cdot F_{i}^{S}$
15      $F_{i}^{\mathcal{I}CS}=Attn(\mathbf{Q}^{\prime}_{i},\mathbf{K}^{\prime}_{i},\mathbf{V}^{\prime}_{i})$
16      # Add and Normalization.
17      $\tilde{F}_{i}^{\mathcal{I}CS}=\text{Norm}(F_{i}^{\mathcal{I}}+\bm{\gamma}\>\text{Dropout}(F_{i}^{\mathcal{I}CS}))$
18      # Self Attention. $\tilde{\mathbf{Q}}_{i},\tilde{\mathbf{K}}_{i},\tilde{\mathbf{V}}_{i}=\tilde{W_{q}}\cdot\tilde{F}_{i}^{\mathcal{I}CS},\tilde{W_{k}}\cdot\tilde{F}_{i}^{\mathcal{I}CS},\tilde{W_{v}}\cdot\tilde{F}_{i}^{\mathcal{I}CS}$
19      $\tilde{F}_{i}^{\mathcal{I}}=Attn(\tilde{\mathbf{Q}}_{i},\tilde{\mathbf{K}}_{i},\tilde{\mathbf{V}}_{i})$
20 end for
21# 3D Scene Reconstruction and Novel View Synthesis.
22Predict $J$ Gaussians $\{(\bm{\mu}_{j},\bm{\alpha}_{j},\bm{\Sigma}_{j},\bm{c}_{j})\}^{J}_{j}$ from $\tilde{F}_{i}^{\mathcal{I}}$.
Render ^$\hat{\mathcal{I}}_{t}$ images with rasterization function.
Algorithm 1 3D scene from a single-view image.
8 Experimental Setup
8.1 Datasets
RealEstate10K. The RealEstate10K [65] dataset consists of large-scale home walkthrough videos from YouTube, including approximately 10 million frames from around 80,000 videos. It also provides camera parameters for each frame calibrated using the Structure-from-Motion (SfM) software. We follow the standard training and testing split, with 67,477 scenes for training and 7,289 for evaluation.
NYUv2. The NYUv2 [43] dataset provides video sequences from diverse indoor environments captured using Kinect cameras. In line with [45], we employ 250 source images from 80 scenes for cross-dataset evaluation and randomly sample target frames within a $\pm$30 frame range from the source, following the random protocol of RE10K [65]. For camera trajectories, we use SfM software as RE10K.
ACID. The ACID [28] dataset consists of large-scale natural landscape videos captured by aerial drones. Like the RE10K [65], ACID provides camera parameters for frames, which are calculated via SfM software. For cross-dataset evaluation, we utilize 450 source images from 150 scenes and randomly sample target frames within a $\pm$30 frame range from the source as the random protocol of RE10K. Note that we evaluate and visualize Flash3D [45] on ACID using publicly available code and provided checkpoints.
KITTI. The KITTI [17] is a landmark autonomous driving dataset containing 30 city driving sequences. Following the well-established evaluation protocol from Tulsiani et al. [50], we utilize 1,079 source frames and provided corresponding camera parameters for cross-dataset evaluation.
8.2 Implementation Details
Our experimental setup is built on the prevalent deep learning framework, PyTorch. For image processing, we use the ResNet-50 [18] image encoder and the UniDepth [39] pre-trained model for monocular depth estimation, with a single image size of $256\times 384$. We employ LLaVA [30] 13B for text embeddings and extend the PointNet [40] encoder for extracting point features. Note that we precompute text embeddings to optimize training efficiency by minimizing computational overhead. Our multi-resolution transformer comprises three layers with 8-headed attention, leveraging three different resolution image features to effectively capture both global structures and fine details. We also set the ratio $\gamma$ as 0.5 to strike a balance, preventing excessive loss of core visual information from image features while integrating our two novel priors. Then, our Gaussian decoder predicts two sets of depth offsets and 3D offsets for vivid scene representation. We use a single A100 GPU for training and select the best-performing model after convergence. Specifically, we optimize a combination of $\mathcal{L}_{\ell 1}$, $\mathcal{L}_{\textrm{ssim}}$, and $\mathcal{L}_{\textrm{lpips}}$ losses using the Adam optimizer with each coefficient as $\lambda_{\ell 1}$=$1$, $\lambda_{\text{ssim}}$=$0.85$, and $\lambda_{\text{lpips}}$=$0.01$, respectively. We will also make the code publicly available for further research.
9 Additional Experiments
9.1 Ablation Studies in Cross-dataset Settings
Method $n=\textit{Random}$ (frames)
Baseline Contextual Spatial PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
$\checkmark$ - - 25.11 0.775 0.178
$\checkmark$ $\checkmark$ - 25.51 0.779 0.163
$\checkmark$ - $\checkmark$ 25.48 0.778 0.165
$\checkmark$ $\checkmark$ $\checkmark$ 25.57 0.781 0.157
Table 8: Ablation study to see the effect of our two priors on the NYUv2 [43] in cross-dataset settings. The “Baseline” refers to our basic transformer architecture without any proposed priors.
Method $n=\textit{Random}$ (frames)
Baseline Contextual Spatial PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
$\checkmark$ - - 24.26 0.732 0.261
$\checkmark$ $\checkmark$ - 24.57 0.735 0.253
$\checkmark$ - $\checkmark$ 24.62 0.737 0.254
$\checkmark$ $\checkmark$ $\checkmark$ 24.73 0.739 0.250
Table 9: Ablation study to see the effect of our two priors on the ACID [65] dataset in cross-dataset settings. The “Baseline” refers to our basic transformer architecture without any proposed priors.
Figure 11: Ablation study to see the effect of iteratively incorporating our novel priors on the RE10K [65] ($n$=Random). For clear ablations, we keep the number of entire transformer layers consistent across the experiments and adjust only the number of cross-attentions (CA).
In this section, we validate the effectiveness of our two innovative priors through ablative experiments across cross-dataset settings. In Tab. 8 and Tab. 9, we evaluate variants of our method, with/ and w/o Contextual and Spatial priors, on the NYUv2 [43] and ACID [28] datasets, respectively. As repeatedly mentioned in the main paper, the Baseline denotes our basic transformer architecture, excluding cross-attention with any of our proposed priors.
First, incorporating contextual cues leads to significant improvements, both for indoor scenes (NYUv2) and outdoor nature scenes (ACID). With text embeddings from a well-trained visual-langualge model (VLM) [30], our network learns not just basic object types or scene semantics but also deeper context, such as how objects relate to each other or the overall structure of the scene. In other words, we take advantage of text embeddings to provide comprehensive general knowledge as well as scene-specific details for generalizable scene reconstruction across diverse environments. Then, these backgrounds serve as effective guidance to capture helpful cues even from the text embeddings of unfamiliar scenes, reconstructing robust 3D scenes.
Additionally, by incorporating spatial guidance, our approach boosts generalization performance on both datasets. Beyond the geometric cues from 2D depth maps, we guide our network to be aware of three-dimensional domains, more associated with 3D Gaussians, through 3D point features. Based on deep spatial understandings, our network effectively reconstructs 3D scenes with accurate Gaussians, even in complex, unfamiliar environments. Finally, combining all priors together achieves further advances, seamlessly complementing limited knowledge from a single-view image. In addition to Tab.4 in our main paper, these results demonstrate the significance of our two novel priors.
9.2 Iteratively Incorporating Priors
In addition to Fig.5 in the main paper, we present additional ablative experimental results to highlight the benefits of iteratively incorporating our priors in Fig. 11. Consistent with the settings in Fig.5 (main), we randomly sample the target frame within a $\pm$30 range; also, fix the total number of transformer layers at three and apply cross-attention either in the first layer only, across two layers, or throughout all three layers. Through iterative cross-attention between image features and our priors, blurry artifacts gradually fade, sharpening the object contours and enhancing clarity in images. Simultaneously, errors between rendered images and target images also steadily decrease. In essence, iterative incorporations of valuable knowledge from our novel priors lead to noticeable improvements in overall visual quality. These findings emphasize both the importance of our priors and the structural robustness of our transformer architecture for challenging monocular 3D scene reconstruction.
9.3 Discussion on Text Descriptions
For rich contextual cues, we leverage text embeddings from a well-trained VLM [30]. Specifically, we prompt the VLM to generate text descriptions for the input image; then, we utilize intermediate text embeddings before they are processed into linguistic description outputs. To discover the optimal text embeddings for 3D scene reconstruction, we investigate the impact of contextual information within various types of text embeddings on generalizability, as shown in Tab.5 of our main paper. For comparison, we conduct experiments with four different styles of prompts: identifying the scene type, listing objects, describing the scene with a detailed single sentence, and two or more sentences. We provide examples of text description outputs using these prompts in Fig. 12. Usually, a single sentence captures comprehensive details for the scene, including textures (e.g., “wooden”, “leather”), object relationships (e.g., “on the countertop”, “surrounded by chairs”, “large mirror above it”), and overall composition (e.g., “on the left side”, “on the outside”), surpassing simple cues like scene type or object list. However, extended sentences often introduce exaggerated or fabricated elements, such as overly interpretive moods, atmospheric descriptions with excessive adjectives (e.g., “organized and inviting”, “adding an artistic touch”), or entirely false specifics (e.g., “two people are present inside the home…”, “lucky numbers…”). These noisy overstatements hinder the network from learning meaningful context information of the text embeddings, resulting in relatively lower performance than using a single sentence. Ultimately, in this work, we benefit from employing well-crafted single sentences to enhance image features with valuable contextual cues, achieving context-aware 3D scene reconstruction with superior novel view synthesis.
Figure 12: Examples of four different formats of text descriptions from the VLM [30], as described in Tab.5 in the main paper.
9.4 Text Embeddings from Various VLMs
Contextual cues from text embeddings are one of our core methods to break through the inherent constraints in monocular settings. Thus, identifying the most effective text embeddings is crucial for achieving high-quality single-view 3D scene reconstruction. In Tab. 10, we explore how text embeddings from various latest pre-trained VLMs, including OpenFlamingo [5], BLIP2 [27] T5, LLaVA [30] 7B, and LLaVA 13B, influence performance on the RE10K [65] dataset. For a fair comparison, we prompt all VLM to produce a single sentence description for the scene. Then, we utilize intermediate text embeddings from each VLM. Even with similar prompts, each model generates distinct structures of text descriptions. For example, OpenFlamingo tends to produce relatively unstable text descriptions with redundant or exaggerated information, providing limited value for 3D scene reconstruction. Meanwhile, BLIP2 and LLaVA 7B generate monotonous text descriptions that primarily focus on object and scene types. On the other hand, LLaVA 13B yields more informative text descriptions with useful details for 3D scene reconstruction, such as textures (e.g., “wooden”, “leather”), object relationships (e.g., “on the countertop”, “surrounded by chairs”, “large mirror above it”), and scene composition (e.g., “on the left side”, “on the outside”), as shown in Fig. 12. Ultimately, we leverage text embeddings from the well-aligned multimodal space of LLaVA 13B, trained on large-scale real-world data, towards context-aware 3D scene reconstruction, going beyond the limited visual cues from a single-view image.
$n=\text{10}$ (frames) $n=\textit{Random}$ (frames)
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
OpenFlamingo 26.08 0.858 0.131 25.06 0.832 0.158
BLIP2 T5 26.29 0.860 0.129 25.27 0.833 0.156
LLaVA 7B 26.19 0.861 0.129 25.23 0.834 0.156
LLaVA 13B 26.40 0.864 0.127 25.40 0.838 0.153
Table 10: Ablation study to see the impact of text features from various VLMs, including OpenFlamingo [5], BLIP2 [27], and LLaVA [30], on 3D scene reconstruction using the RE10K [65].
9.5 Visual Comparison
We present additional qualitative comparisons across the RE10K [65] in Fig. 14 and Fig. 15 as well as ACID [28] (Fig. 16) and KITTI [17] (Fig. 17) in cross-dataset settings.
10 Limitations and Future Work
Figure 13: Failure cases of CATSplat. When invisible areas in the input become visible in the target, ours might be less productive.
Although CATSplat shines in monocular 3D scene reconstruction with two additional priors, it does not ensure perfect novel view synthesis across all real-world scenarios. Depending on dynamic camera movements, when regions that are occluded, truncated, or even entirely missing in the input image appear in the target view, ours might be less effective. For example, in Fig. 13, when previously unseen elements, like green plants absent in the input, emerge in the target view (Scene1) or when areas of the bathroom, once hidden behind a door, become visible (Scene2), our model struggles to reconstruct these newly revealed parts. In the future, we plan to explore involving generative knowledge to better handle these unseen regions in monocular 3D scene reconstruction. Moreover, we believe that training the model on a broader range of datasets will strengthen its general understanding of challenging natural environments.
Figure 14: Qualitative comparisons between Flash3D [45] and Ours with Input Image and Ground Truth on the RealEstate10K [65] dataset.
Figure 15: Qualitative comparisons between Flash3D [45] and Ours with Input Image and Ground Truth on the RealEstate10K [65] dataset.
Figure 16: Qualitative comparisons between Flash3D [45] and Ours with Input Image and Ground Truth on the ACID [28] dataset.
Figure 17: Qualitative comparisons between Flash3D [45] and Ours with Input Image and Ground Truth on the KITTI [17] dataset.